# Fair Conformal Prediction for Incomplete Covariate Data

### v1

@@ Abstract
Conformal prediction provides a distribution-free framework for uncertainty quantification. This study explores the application of conformal prediction in scenarios where covariates are missing, which introduces significant challenges for uncertainty quantification. We establish that marginal validity holds for imputed datasets across various mechanisms of missing data and most imputation methods. Building on the framework of nonexchangeable conformal prediction, we demonstrate that coverage guarantees depend on the mask. To address this, we propose a nonexchangeable conformal prediction method for missing covariates that satisfies both marginal and mask-conditional validity. However, as this method does not ensure asymptotic conditional validity, we further introduce a localized conformal prediction approach that employs a novel score function based on kernel smoothing. This method achieves marginal, mask-conditional, and asymptotic conditional validity under certain assumptions. Extensive simulation studies and real-data analysis demonstrate the advantages of these proposed methods.
@@ 1 Introduction
Statistical and machine learning algorithms are increasingly important for high-stakes decision-making processes in various fields, including autonomous driving, healthcare, and civic applications. As the use of complex black-box methods, like deep neural networks, grows, so does the cost of making wrong decisions, underscoring the need for robust methods to quantify predictive uncertainty. Recent years have seen a rise in interest regarding uncertainty quantification techniques, especially those based on conformal prediction. See e.g. [37, 34, 22, 36, 1, 4]. Conformal prediction aims to build a prediction set for a response variable $Y\in\mathcal{Y}$ after observing a random vector $X\in\mathcal{X}\subseteq\mathbb{R}^{d}$ that contains $d$-dimensional explanatory variables. By setting a specific miscoverage level $\alpha\in(0,1)$, conformal prediction creates a prediction set $\mathcal{C}_{\alpha}(\cdot)$ that meets the following coverage guarantee:
@@
This approach provides a finite-sample coverage guarantee for any underlying predictive algorithm, such as neural networks or random forests, under fairly mild conditions. As a result, conformal prediction is suitable for a wide range of real-world applications, delivering both flexibility and reliability in uncertainty quantification.
@@
From a fairness perspective, each subgroup defined by a specific mask should receive a prediction interval with equal probability of containing the true response. Moreover, missingness patterns may correlate with unobserved external features linked to meaningful categories. For example, in survey data, affluent individuals are often less likely to disclose their income, leading to missing income variables. As a result, these masks may relate to socio-economic information.
@@
The approach to handling missing data depends on the specific downstream task. This topic has been extensively studied in various statistical contexts, as noted in works by [15] and [24]. Recent works study missing data in several different areas, including principal component analysis [43], U-statistics [6], changepoint detection [8], testing whether the missingness is independent of the data [5] and classification [32]. Moreover, it also receives a lots of attention in the framework of prediction [17, 18, 19, 2, 3, 40, 41, 16]. For the uncertainty quantification with missing data, [11] and [35] consider uncertainty quantification for matrix completion tasks; [31] consider the problem of missingness in the response, which is generally known as semi-supervised setting. Under the same framework, [20] leverages tools from causal inference to achieve strong guarantees such as feature and outcome’s missingness conditional coverage.
@@
Our work on uncertainty quantification with missing data builds builds upon the foundational work of [40] and [41], which examine the predictive framework of estimating $Y$ given $X$, where $X$ may exhibit missing values during both training and testing phases. In [40], two conformal methods were proposed, named CQR-MDA-Exact and CQR-MDA-Nested, both of which achieve marginal and mask-conditional validity (defined in (1) and (2) below) under specific conditions. Building on this, [41] further provided a comprehensive framework in which CQR-MDA-Exact and CQR-MDA-Nested are established as special cases. However, the theoretical implications of missingness on marginal and mask-conditional validity remain somewhat ambiguous. Additionally, the assumption related to the missingness mechanism is a bit restrictive. Note that [32] recently proposed a general missingness mechanism specifically for binary classification tasks. Motivated by [32], we adapt this mechanism for the regression context and build upon the concept of nonexchangeable conformal prediction, as laid out by [4], to investigate the influence of missingness on both marginal and mask-conditional validity.
@@
In this article, we consider conformal prediction in the presence of missing values within the covariates. Our focus is on understanding how such missingness affects the validity and efficiency of conformal prediction. To effectively address the challenges posed by missing covariates, we introduce two novel conformal methods. The key contributions of this work are summarized as follows: (1) We show that conformal prediction is marginally valid regardless of the model, missing data mechanism, and imputation function. However, the coverage probabilities associated with various missing data patterns differ, and we theoretically prove that the coverage probability is ultimately constrained by a summation of total variance. (2) We introduce a nonexchangeable conformal prediction method for missing covariates. This approach selects applicable calibration cases and assigns weights based on a distance metric that reflects similarity in missingness. We prove that the method is both marginally and mask-conditionally valid. (3) While the nonexchangeable method does not guarantee asymptotic conditional validity, we further propose a localized conformal method using kernel smoothing and a refined score function. This method is shown to be marginally valid, mask-conditionally valid, and asymptotically conditionally valid under certain assumptions.
@@ 2 Preliminaries
[…]
@@
Marginal validity (1) indicates that the coverage guarantee is satisfied for all data across different masks as a collective group. Although marginal validity is a desirable property of a prediction interval, achieving mask-conditional validity in a distribution-free and finite-sample context is equally important. The masks create meaningful categories, and ensuring mask-conditional validity (2) promotes more equitable treatment [40].
Since most statistical models and machine learning methods are not inherently designed to handle incomplete data, a common approach to utilizing existing pipelines in the presence of missing values is to impute the data, thereby creating a complete dataset. We will utilize an imputation function that maps observed values to themselves while assigning plausible values to the missing entries based on those observed values. Following the notations defined in [19], we define the imputation function $\phi^{M}:\mathbb{R}^{|\text{obs}(M)|}\rightarrow\mathbb{R}^{|\text{mis}(M)|}$, where the input consists of the observed values and the output comprises the imputed values, conditional on a mask $M\in\{0,1\}^{d}$. The imputation function belongs to the function class defined as follows:
@@
Symmetrical condition is a mild condition relative to the regression algorithm [37, 22, 4]. In addition, most existing imputation methods for exchangeable (a condition weaker than i.i.d.) data satisfy the symmetrical condition [40].
Theorem 1 (Marginal validity).
@@
By the i.i.d. and symmetrical algorithms assumptions, the scores are exchangeable. This result shows that marginal validity holds regardless of the missing mechanism (MCAR, MAR or MNAR) and the symmetric imputation scheme. We have shown that the prediction interval (6) is marginally valid (1), under mild conditions; however, achieving mask-conditional validity (2) is a challenging task. To better illustrate the interplay between missing values and uncertainty quantification with respect to the mask, we consider the following example.
Example 1 (Gaussian linear model).
Suppose that the data is generated according to a linear model
$Y=\beta^{T}X+\varepsilon,$
where $\varepsilon\sim N(0,\sigma^{2})\perp\!\!\!\perp(X,M),\beta\in\mathbb{R}^{d}$, and the covariates are Gaussian pattern mixture model [23], i.e., for all $m\in\{0,1\}^{d}$, there exist $\mu^{m}$ and $\Sigma^{m}$ such that $X|(M=m)\sim N(\mu^{m},\Sigma^{m})$. Without loss of generality, let $X=(X_{1},X_{2})^{T}$, where $X_{1}$ is missing, $X_{2}$ is observed, and their coefficients $\beta=(\beta_{1},\beta_{2})^{T}$. The Gaussian pattern mixture model is denoted as
$(X_{1},X_{2})^{T}\mid M=m\sim N\left(\begin{pmatrix}\mu_{1}^{m}\\ \mu_{2}^{m}\end{pmatrix},\begin{pmatrix}\Sigma_{11}^{m}&\Sigma_{12}^{m}\\ \Sigma_{21}^{m}&\Sigma_{22}^{m}\end{pmatrix}\right).$ (7)
Then, the distribution of $Y$ given $X_{\text{obs}(m)},M=m$ is
$Y\mid(X_{\text{obs}(m)},M=m)\sim N\left(\beta_{2}^{T}X_{2}+\beta_{1}^{T}\mu_{1\mid 2}^{m},\beta_{1}^{T}\Sigma_{1\mid 2}^{m}\beta_{1}+\sigma^{2}\right),$ (8)
where $\mu_{1\mid 2}^{m}=\mu_{1}^{m}+\Sigma_{12}^{m}{\Sigma_{22}^{m}}^{-1}(X_{2}-\mu_{2})$ and $\Sigma_{1\mid 2}^{m}=\Sigma_{11}^{m}-\Sigma_{12}^{m}{\Sigma_{22}^{m}}^{-1}\Sigma_{21}^{m}$.
The result is straightforward using the property of the conditional distribution of multivariate normal distribution, and the calculation is provided in the appendix. The Gaussian pattern model considered in (7) can be encoded for different scenarios including MCAR, MAR, or MNAR settings [2]. Recall that our goal is to build a prediction interval for $Y$ given $X_{\text{obs}(m)}$ and $M=m$. Equation (8) emphasizes that even when the noise of the generative model is homoscedastic, missing values induce heteroskedasticity. Indeed, the variance of the conditional distribution of $Y\mid(X_{\text{obs}(m)},M=m)$ depends on missing pattern $m$. Further, the uncertainty increases as missing values are associated with larger regression coefficients.
@@ 3.1 Nonexchangeable conformal prediction
which means that the interval length remains invariant regardless of the mask. However, equation (8) shows that the prediction interval length with respect to different masks should be variant. As missing values induce heteroskedasticity, to build a prediction interval for $Y^{(n+1)}$ given $X_{\text{obs}(m)}^{(n+1)}$ and $M^{(n+1)}=m$, a natural choice is to upweight the data with missingness similar to the test point and downweight those with missingness not similar to the test point.
@@
[…]
@@
Beginning with nonexchangeable conformal prediction, we demonstrate that achieving marginal validity is feasible. However, attaining mask-conditional validity presents considerable challenges. As outlined in Theorem 2, a simplistic approach to achieve mask-conditional validity would be to exclude data points indexed by $i$ such that $M^{(i)}\neq M^{(n+1)}$, retaining only those data points that share the same mask as the test point. This strategy, however, proves unrealistic in scenarios characterized by missing covariates. To illustrate this, consider the missing pattern $m\in\{0,1\}^{d}$. The number of distinct missing patterns tends to grow exponentially with the dimensionality $d$; in fact, there are $12^{d}-$ 1 possible missing patterns. Consequently, the quantity of data available for any specific missing pattern may be exceedingly limited. When reassessing the conditional probability, a more reasonable strategy for achieving mask-conditional validity is to retain some data points where the total variance can be made zero under certain mild conditions, while discarding others that cannot satisfy this requirement. To support this approach, we will introduce some general assumptions about the missingness mechanism.
@@
We say that $(X,M,Y)$ is an available case for $m$, if $M\preceq m$. Then the available cases for the mask of the test point $\left(X^{(n+1)},M^{(n+1)},Y^{(n+1)}\right)$ can be indexed by ~=$\tilde{\mathcal{I}}=\{i\in[n]:M^{(i)}\preceq M^{(n+1)}\}$. Under Assumption 2, the conditional distribution of $Y\mid\left(X_{\text{obs}(M^{(n+1)})},M^{(i)}\right)$, for $i\in\tilde{\mathcal{I}}$, will be equal to that of $Y\mid\left(X_{\text{obs}(M^{(n+1)})},M^{(n+1)}\right)$. Algorithm 1 presents a new approach to nonexchangeable conformal prediction.
@@
5:Find available cases ~=$\tilde{\mathcal{I}}=\left\{i\in[n]:M^{(i)}\preceq M^{(n+1)}\right\}$
@@
8:Rearrange the scores based on the ascending order of the distances $d^{(i)},i\in\tilde{\mathcal{I}}$. The scores are denoted as $\tilde{V}_{*}^{(i)},i\in[|\tilde{\mathcal{I}}|]$
9:Calculate $w_{i}=\begin{cases}\rho^{\left|\tilde{\mathcal{I}}\right|+1-i},&i\in[|\tilde{\mathcal{I}}|]\text{ and }M^{(i)}\neq M^{(n+1)},\\ 1,&i\in[|\tilde{\mathcal{I}}|]\text{ and }M^{(i)}=M^{(n+1)}.\end{cases}$, and $w_{n+1}=1$. Normalize the weights: $\tilde{w}_{i}=w_{i}/\left(w_{n+1}+{\textstyle\sum_{j\in[|\tilde{\mathcal{I}}|]}}w_{j}\right),i\in[|\tilde{\mathcal{I}}|]$, and $\tilde{w}_{n+1}=w_{n+1}/\left(w_{n+1}+{\textstyle\sum_{j\in[|\tilde{\mathcal{I}}|]}}w_{j}\right)$.

### v2

@@ Abstract
Conformal prediction provides a distribution-free framework for uncertainty quantification. This study explores the application of conformal prediction in scenarios where covariates are missing, which introduces significant challenges for uncertainty quantification. Building on the framework of nonexchangeable conformal prediction, we demonstrate that coverage guarantees depend on the mask. To address this, we propose a nonexchangeable conformal prediction method for missing covariates that satisfies both marginal and mask-conditional validity. However, as this method does not ensure asymptotic conditional validity, we further introduce a localized conformal prediction approach that employs a novel score function based on kernel smoothing. This method achieves marginal, mask-conditional, and asymptotic conditional validity under certain assumptions. Extensive simulation studies and real-data analysis demonstrate the advantages of these proposed methods.
@@ 1 Introduction
Conformal prediction is a general framework for distribution-free uncertainty quantification [41, 38, 26, 40, 1, 6]. Given a random vector $X\in\mathcal{X}\subseteq\mathbb{R}^{d}$, which contains $d$-dimensional explanatory variables, the goal of conformal prediction is to construct a prediction set $\mathcal{C}_{\alpha}(\cdot)$ for the response variable $Y\in\mathcal{Y}$, such that for a pre-specified miscoverage level $\alpha\in(0,1)$, the following coverage guarantee holds:
@@
This guarantee holds under minimal assumptions and for any underlying predictive model, including complex black-box algorithms. Consequently, conformal prediction has gained traction in a wide range of real-world applications due to its flexibility, robustness, and finite-sample validity. Notable applications include survival data analysis [9], causal inference [44], outlier detection [7], risk control [2], time series forecasting [1], and selective inference [5], among others.
@@
From a fairness perspective, each subgroup defined by a specific mask should receive a prediction interval with equal probability of containing the true response. Moreover, missingness patterns may correlate with unobserved external features linked to meaningful categories. For example, in survey data, affluent individuals are often less likely to disclose their income, leading to missing income variables. As a result, these masks may relate to socio-economic information [45].
@@
The approach to handling missing data depends on the specific downstream task. This topic has been extensively studied in various statistical contexts [28]. Recent works study missing data in several different areas, including principal component analysis [48], U-statistics [10], changepoint detection [12], testing whether the missingness is independent of the data [8] and classification [36]. Moreover, it also receives a lots of attention in the framework of prediction [21, 22, 23, 3, 4, 15, 39, 35, 45, 46, 20].
@@
Our work on uncertainty quantification with missing data builds builds upon the foundational work of [45, 46], which examine the predictive framework of estimating $Y$ given $X$, where $X$ may exhibit missing values during both training and testing phases. In [45], two conformal methods were proposed, named CQR-MDA-Exact and CQR-MDA-Nested, both of which achieve marginal and mask-conditional validity (defined in (1) and (2) below) under specific conditions. Building on this, [46] further provided a comprehensive framework in which CQR-MDA-Exact and CQR-MDA-Nested are established as special cases. However, the theoretical implications of missingness on marginal and mask-conditional validity remain somewhat ambiguous. Additionally, the assumption related to the missingness mechanism is a bit restrictive. Note that [36] recently proposed a general missingness mechanism specifically for binary classification tasks. Motivated by [36], we adapt this mechanism for the regression context and build upon the concept of nonexchangeable conformal prediction, as laid out by [6], to investigate the influence of missingness on both marginal and mask-conditional validity.
@@
In this article, we consider conformal prediction in the presence of missing values within the covariates. Our focus is on understanding how such missingness affects the validity and efficiency of conformal prediction. To effectively address the challenges posed by missing covariates, we introduce two novel conformal methods. The key contributions of this work are summarized as follows: (1) We show that the coverage probabilities associated with various missing data patterns differ and theoretically establish that the coverage probability is ultimately constrained by a summation of total variance. (2) We introduce a nonexchangeable conformal prediction method for missing covariates. This approach selects applicable calibration cases and assigns weights based on a distance metric that reflects similarity in missingness. We prove that the method is both marginally and mask-conditionally valid. (3) While the nonexchangeable method does not guarantee asymptotic conditional validity, we further propose a localized conformal method using kernel smoothing and a refined score function. This method is shown to be marginally valid, mask-conditionally valid, and asymptotically conditionally valid under certain assumptions.
@@ 2 Preliminaries
For convenience, some notations are also introduced. For $n\in\mathbb{N}$, let $n]=\{1,\cdots,n\}$. Denote $|\mathcal{I}|$ as the cardinality of the index set $\mathcal{I}$. For $\tilde{m}\in\{0,1\}^{d}, mm~$$\preceq\tilde{m}$ means that $m_{i}\leq\tilde{m}_{i},i\in[d]$, where $m_{i}$ (resp. $\tilde{m}_{i}$) is the $i$th component of the vector $m$ (resp. $\tilde{m}$). For two random variables $X$ and $Y$, $X\perp\!\!\!\perp Y$ means that $X$ and $Y$ are independent, and $X\overset{\text{d}}{=}Y$ means that $X$ and $Y$ have the same distribution.
@@
Marginal validity (1) indicates that the coverage guarantee is satisfied for all data across different masks as a collective group. Although marginal validity is a desirable property of a prediction interval, achieving mask-conditional validity in a distribution-free and finite-sample context is equally important. The masks create meaningful categories, and ensuring mask-conditional validity (2) promotes more equitable treatment [45, 46].
Since most statistical models and machine learning methods are not inherently designed to handle incomplete data, a common approach to utilizing existing pipelines in the presence of missing values is to impute the data, thereby creating a complete dataset. We will utilize an imputation function that maps observed values to themselves while assigning plausible values to the missing entries based on those observed values. Following the notations defined in [45, 46], we define the imputation function $\phi^{M}:\mathbb{R}^{|\text{obs}(M)|}\rightarrow\mathbb{R}^{|\text{mis}(M)|}$, where the input consists of the observed values and the output comprises the imputed values, conditional on a mask $M\in\{0,1\}^{d}$. The imputation function belongs to the function class defined as follows:
@@
Symmetrical condition is a mild condition relative to the regression algorithm [41, 26, 6]. In addition, most existing imputation methods for exchangeable (a condition weaker than i.i.d.) data satisfy the symmetrical condition [45, 46].
Theorem 1 (Marginal validity [45, 46]).
@@
By the i.i.d. and symmetrical algorithms assumptions, the scores are exchangeable. This result shows that marginal validity holds regardless of the missing mechanism (MCAR, MAR or MNAR) and the symmetric imputation scheme. The prediction interval (6) is marginally valid (1), under mild conditions; however, achieving mask-conditional validity (2) is a challenging task. To better illustrate the interplay between missing values and uncertainty quantification with respect to the mask, we consider the following example.
Example 1.
Consider the following data-generating process:
$Y=\beta_{1}X_{1}+\beta_{2}X_{2}+\varepsilon,$
where $\varepsilon\sim N(0,\sigma^{2})$ is independent of both $X$ and $M$, $\beta_{1},\beta_{2}$ are scalars, and
$\begin{pmatrix}X_{1}\\ X_{2}\end{pmatrix}\sim N\left(\begin{pmatrix}\mu_{1}\\ \mu_{2}\end{pmatrix},\begin{pmatrix}\sigma_{1}^{2}&\rho\sigma_{1}\sigma_{2}\\ \rho\sigma_{1}\sigma_{2}&\sigma_{2}^{2}\end{pmatrix}\right)$.
The missing data mechanism is defined as follows: $M_{1}=1$ if $X_{1}>\tau_{1}$, and $M_{2}=1$ if $X_{2}<\tau_{2}$. Then, the conditional variance of the response variable given the observed covariates and a specific mask $m$ is
$\text{Var}(Y\mid X_{\text{obs}(m)},M=m)=\begin{cases}\sigma^{2},&m=(0,0)\\[6.0pt] \beta_{1}^{2}\sigma_{1|2}^{2}\left(1+\dfrac{\frac{\tau_{1}-\mu_{1|2}}{\sigma_{1|2}}\,\phi\left(\frac{\tau_{1}-\mu_{1|2}}{\sigma_{1|2}}\right)}{1-\Phi\left(\frac{\tau_{1}-\mu_{1|2}}{\sigma_{1|2}}\right)}-\left[\dfrac{\phi\left(\frac{\tau_{1}-\mu_{1|2}}{\sigma_{1|2}}\right)}{1-\Phi\left(\frac{\tau_{1}-\mu_{1|2}}{\sigma_{1|2}}\right)}\right]^{2}\right)+\sigma^{2},&m=(1,0)\\[10.0pt] \beta_{2}^{2}\sigma_{2|1}^{2}\left(1-\dfrac{\frac{\tau_{2}-\mu_{2|1}}{\sigma_{2|1}}\,\phi\left(\frac{\tau_{2}-\mu_{2|1}}{\sigma_{2|1}}\right)}{\Phi\left(\frac{\tau_{2}-\mu_{2|1}}{\sigma_{2|1}}\right)}-\left[\dfrac{\phi\left(\frac{\tau_{2}-\mu_{2|1}}{\sigma_{2|1}}\right)}{\Phi\left(\frac{\tau_{2}-\mu_{2|1}}{\sigma_{2|1}}\right)}\right]^{2}\right)+\sigma^{2},&m=(0,1)\end{cases}$
where
$\mu_{1|2}=\mu_{1}+\rho\frac{\sigma_{1}}{\sigma_{2}}(X_{2}-\mu_{2}),\quad\sigma_{1|2}^{2}=\sigma_{1}^{2}(1-\rho^{2}),\quad\mu_{2|1}=\mu_{2}+\rho\frac{\sigma_{2}}{\sigma_{1}}(X_{1}-\mu_{1}),\quad\sigma_{2|1}^{2}=\sigma_{2}^{2}(1-\rho^{2}),$
and $\phi(\cdot)$ and $\Phi(\cdot)$ denote the probability density function and cumulative distribution function of the standard normal distribution, respectively.
Example 1 emphasizes that even when the noise of the generative model is homoscedastic, missing values induce heteroskedasticity. Indeed, the variance of the conditional distribution of $Y\mid(X_{\text{obs}(m)},M=m)$ depends on missing pattern $m$.
@@ 3.1 Nonexchangeable conformal prediction
which means that the interval length remains invariant regardless of the mask. However, Example 1 shows that the prediction interval length with respect to different masks should be variant. As missing values induce heteroskedasticity, to build a prediction interval for $Y^{(n+1)}$ given $X_{\text{obs}(m)}^{(n+1)}$ and $M^{(n+1)}=m$, a natural choice is to upweight the data with missingness similar to the test point and downweight those with missingness not similar to the test point.
@@
In the context of the marginal validity of the theorem, the data points $(X^{(i)},M^{(i)},Y^{(i)})$ for $i\in[n+1]$ are indeed independent and identically distributed (i.i.d.). Consequently, the associated scores $V^{(i)}$ for $i\in[n+1]$ exhibit exchangeability. This leads to the equivalence in distribution $R\overset{\text{d}}{=}R^{i}$, and it follows that the total variation distance $\text{d}_{\text{TV}}(R,R^{i})=0$ holds for all $i$. Importantly, the introduction of fixed weights $w_{i}$ into the quantile calculations for constructing the prediction interval does not result in any loss of coverage with respect to marginal validity.
@@
Beginning with nonexchangeable conformal prediction, we demonstrate that achieving marginal validity is feasible. However, attaining mask-conditional validity presents considerable challenges. As outlined in Theorem 2, a simplistic approach to achieve mask-conditional validity would be to exclude data points indexed by $i$ such that $M^{(i)}\neq M^{(n+1)}$, retaining only those data points that share the same mask as the test point. This strategy, however, proves unrealistic in scenarios characterized by missing covariates. To illustrate this, consider the missing pattern $m\in\{0,1\}^{d}$. The number of distinct missing patterns tends to grow exponentially with the dimensionality $d$; in fact, there are $2^{d}-1$ possible missing patterns. Consequently, the quantity of data available for any specific missing pattern may be exceedingly limited [45, 46]. When reassessing the conditional probability, a more reasonable strategy for achieving mask-conditional validity is to retain some data points where the total variance can be made zero under certain mild conditions, while discarding others that cannot satisfy this requirement. To support this approach, we will introduce some general assumptions about the missingness mechanism.
@@
We say that $(X,M,Y)$ is an available case for $m$, if $M\preceq m$. Then the available cases for the mask of the test point $\left(X^{(n+1)},M^{(n+1)},Y^{(n+1)}\right)$ can be indexed by $\tilde{\mathcal{I}}=\{i\in[n]:M^{(i)}\preceq M^{(n+1)}\}$. Under Assumption 2, the conditional distribution of $Y\mid\left(X_{\text{obs}(M^{(n+1)})},M^{(i)}\right)$, for $i\in\tilde{\mathcal{I}}$, will be equal to that of $Y\mid\left(X_{\text{obs}(M^{(n+1)})},M^{(n+1)}\right)$. Algorithm 1 presents a new approach to nonexchangeable conformal prediction.
@@
5:Find available cases $\tilde{\mathcal{I}}=\left\{i\in[n]:M^{(i)}\preceq M^{(n+1)}\right\}$
@@
[…]